Exercise 21-7 On January 1, 2017, Monty Company leased equipment to Flounder Corporation. The following information pertains to this lease. 1. The term of the noncancelable lease is 6 years, with no renewal option. The equipment reverts to the lessor at the termination of the lease 2. Equal rental payments are due on January 1 of each year, beginning in 2017 3. The fair value of the equipment on January 1, 2017, is $127,000, and its cost is $101,600 4. The equipment has an economic life of 8 years, with an unguaranteed residual value of $11,000. Flounder depreciates all of its equipment on a straight-line basis. 5. Monty set the annual rental to ensure an 12% rate of retum. Flounders incremental borrowing rate is 13%, and the implicit rate of the lessor is unknown. 6. Collectibility of lease payments is reasonably predictable, and no important uncertainties surround the amount of costs yet to be incurred by the lessor. Both the lessor and the lessees accounting period ends on December 31.) Calculate the amount of the annual rental payment. (Round present value factor calculations to 5 decimal places, e.g. 1.25124 and the final answer to 0 decimal places e.g. 58,971.) The amount of the annual rental payment Prepare all the necessary journal entries for Flounder for 2017. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, selecit No Entry for the account titles and enter 0 for the amounts. Round present value factor calculations to 5 decimal places, e.g. 1.25124 and the final answer to decimal places e.g. 58,971.) Date Account Titles and Explanation Debit Credit To record the lease.) To record lease paymant.) To record depreciation.)To record interest.) SHOW LIST OF ACCOUNTS Prepare all the necessary journal entries for Monty for 2017. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select No Entry for the account titles and enter 0 for the amounts. Round present value factor calculations to 5 decimal places, e.g. 1.25124 and the final ansver to 0 decimal places e.g. 58,971.) Date Account Titles and Explanation Debit Credit To record the lease.) To record lease payment.) 12/31/17 (To record interest.) Click if you would like to Show Work for this question: Open Show Work LIST OF ACCOUNTS
Answer #1

Part A

The amount of the annual rental payment = $26370

(127000-(11000*0.50663))/ 4.60478 = 26370

Present value of $1 at 12% for 6 periods = 0.50663

Present value of an annuity due at 12% for 6 periods =4.60478

Part B

Date

Account titles and explanation

debit

credit

1/1/17

Leased Equipment

119119

Lease Liability (26370*4.51723)

119119

(To record the lease.)

Lease Liability

26370

Cash

26370

(To record lease payment.)

12/31/17

Depreciation Expense

19853

Accumulated Depreciation (119119/6)

19853

(To record depreciation.)

Interest Expense

12057

Interest Payable (119119-26370)*13%

12057

(To record interest.)

Present value of an annuity due of $1 at 13% for 6 periods = 4.51723

Part C

Date

Account titles and explanation

debit

credit

1/1/17

Lease Receivable

127000

Cost of Goods Sold

96027

Sales Revenue

121427

Inventory

101600

(To record the lease.)

Cash

26370

Lease Receivable

26370

(To record lease payment.)

12/31/17

Interest Receivable

12076

Interest Revenue (127000-26370)*12%

12076

(To record interest.)

Cost of goods sold = 101600-(11000*0.50663) = 96027

Sales revenue = 127000-(11000*0.50663) = 121427